OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b?

OpenStudy (anonymous):

y = −bx − b(1 − a) y = bx − b(a − 1) y=1/bx+b[a-1] y=-1/bx+b[1-a] these are my answer choices